OpenStudy (anonymous):

Write the equation of the line that passes through (3, 6) and (0, –3) in slope-intercept form.

OpenStudy (anonymous):

First find the slope: \[m = \Large \frac{ y2 - y1 }{ x2 - x1 } \to \frac{ -3 - 6 }{ 0 - 3 } \to \frac{ -9 }{ -3 } \to 3\]

OpenStudy (anonymous):

Do you know the Point-Slope form?

OpenStudy (anonymous):

@NasiaNay